# NuiScene: Exploring Efficient Generation of Unbounded Outdoor Scenes

### v1

@@ Abstract
Figure 2: A scene generated by our model, trained on multiple scenes. We texture the scene using SceneTex. As shown, our model combines elements from different scenes in the dataset such as castles and skyscrapers. For the untextured mesh, please see the supplement (Figure 16).
@@ 1 Introduction
Following the trend of diffusion models in image generation—such as Stable Diffusion [36]—generative models for creating standalone 3D objects [45, 23] are also rapidly maturing. The success of such models is partly driven by large-scale datasets like Objaverse [9]. However, large-scale scene generation remains limited, with many works focused on indoor environments [17, 44, 32, 4]. While some methods utilize semantic car datasets [25, 30, 26] or CAD models [29], these datasets lack high-quality geometry and have limited scene variation and height. Compared to previous methods, we tackle a challenging task with three key challenges: 1) enabling fast and efficient generation of outdoor scenes with varying height distributions, 2) generating scenes with varying styles into a cohesive and continuous scene, 3) curating a dataset that facilitates exploration into such methods and supports joint training of heterogeneous scenes.
@@
Previous works on generating unbounded indoor scenes using 3D-Front [11] assume scenes can be divided into equal-sized cubes, and learn spatially structured latents like triplanes [44] or dense feature grids [32] for compression. This approach struggles with outdoor scenes featuring buildings like skyscrapers with significantly different heights. As shown in Figure 3, naively scaling these representations can result in intensive memory usage. Rescaling the scene via normalization will also degrade details. To address this, we compress scene chunks into vector sets using 3DShape2VecSet [49], achieving better compression than triplanes for more efficient training and higher quality generations. Additionally, we train our diffusion model for explicit outpainting, further improving inference speed over methods [44, 32] utilizing resampling-based inpainting like RePaint [31] which requires additional diffusion steps. Furthermore, prior methods focus on uniform datasets, such as indoor or urban driving scenes. In contrast, we aim to explore generative models that can bring together scenes with different context such as castles and cities. We showcase our model’s ability to generate scenes with distinct styles in Figure 2.
Another key limitation for such a task is the lack of openly available high quality outdoor scenes for training, making developing such methods more challenging. 3D-Front [11] is used by many prior works to train unbounded indoor scene generation. However, it lacks variation in scene height and non-furniture geometry. Semantic driving datasets [42, 2] and urban reconstructions [35, 13, 12] lack high quality geometry compared to artist-created meshes. Objaverse [9] includes many scenes that could facilitate training for this task. However, several challenges remain. Notably, these meshes lack a unified global scale, and feature highly varied ground geometries—ranging from very thick to thin planes—making it difficult to train a generative model effectively as a homogeneous dataset. To address this, we curate a small set of scenes from Objaverse. We clean ground geometries and establish a unified scale, enabling the development of methods that handle (1) diverse scene height distributions and (2) outdoor geometry from varied sources and styles. We show that this data process pipeline is effective in enabling the joint training of heterogeneous scenes in a unified generative model.
@@ 2 Related Work
Figure 4: Overview of our VAE and diffusion models. For the VAE we use vector sets [49] for latent compression. For the diffusion model we train a conditional outpainting model with four different settings to allow for fast generation in a raster scan pattern during inference.
@@
3D Bounded Single Exemplar Scene Gen. Several works focus on generating 3D scenes from a single exemplar [41, 18, 43, 27], follow a coarse-to-fine progressive approach by capturing patch statistics across scales. Similar to SinGAN [37], SinGRAV [41], 3INGAN [18], and Wu and Zheng [43] adopt a pyramid of GANs, where discriminators operate on smaller patch sizes to enforce local consistency while allowing diverse global compositions. Li et al. [27] introduces EM-like optimization for patch matching and blending across different scales. Although these methods generate scenes with varying compositions and aspect ratios, they are limited to bounded generation, and producing an entire scene at once leads to increasing memory costs for larger scenes.
3D Unbounded Scene Gen. One common method for 3D unbounded scene generation start by dividing scenes into smaller chunks and using an autoencoder to compress them into latents, such as triplane or dense grids. A diffusion model is then trained on the scene chunk latents. By overlapping the current chunk with previously generated ones, these methods enable continuous outpainting relying on resampling-based inpainting methods like RePaint [31] requiring additional diffusion steps.
@@
There are also other methods that take different approaches to unbounded outdoor generation. CityDreamer [46] and InfiniCity [29] leverage top-down views for unbounded generation but are limited to cityscapes with low-resolution geometry. SceneDreamer [7] and Persistent Nature [5] are promising directions to generate unbounded 3D worlds from only 2D images. However, without explicit 3D supervision, the resulting geometry lacks detail.
For unbounded indoor scene generation, methods typically train on 3D-Front [11] and assume cubic scene chunks. BlockFusion [44], like SemCity, compresses scene chunks into triplanes. It then employs another VAE to further compress the triplanes into a lower resolution, reducing redundancy in the representation to improve diffusion learning. LT3SD [32] encodes TUDF chunks into multi-level dense feature grid latent trees, organizing both its autoencoder and diffusion models in a hierarchical manner for coarse-to-fine chunk generation. While these representations perform well for indoor scenes with fixed ceiling heights, they struggle to scale to outdoor environments with significant height variations, as shown in Figure 3. To tackle this, we use a vector set representation which compresses chunks into a uniform size, offering better compression and performance. And enable faster generation with a outpainting model.
@@ 3 Method
We represent a large outdoor scene as a collection of local scene chunks, where a VAE learns to compress individual chunks into vector sets, and a diffusion model generates neighboring chunks in a $2\times 2$ grid using the VAE’s learned representation. By combining both models, we enable continuous chunk generation, supporting unbounded outdoor scene synthesis. Our overall training framework follows the Latent Diffusion Model (LDM)[36], with an illustration provided in Figure 4.
The dataset curation process is outlined in Sec. 3.1, followed by how training chunks are sampled. To explore efficient large-scale scene generation, we experiment with triplane and vector set latents. The VAE backbone and the process of querying points from these representations to obtain occupancies are described in Sec. 3.2. For fast inference without relying on resampling-based inpainting methods, we train an explicit outpainting diffusion model, detailed in Sec. 3.3. Optionally, textures can be generated using SceneTex [6]. More details are available in the supplement (S2.3).
@@ 3.1 Dataset
Filtering and Preprocessing. We begin by filtering Objaverse [9] using multi-view embeddings from DuoduoCLIP [24] to select 43 high-quality scenes. Next, we establish a consistent scale by labeling scenes with relative scales. We then convert each scene into an occupancy grid with a resolution multiplied by its labeled relative scale and generate meshes using the marching cubes algorithm for point cloud sampling. To ensure uniform ground geometry across scenes, we ensure a uniform thickness below the ground level for each scene. For more details on the dataset curation process, please refer to the supplement (S1).
@@
Chunk Sampling. A sample scene processed through our curation pipeline is shown in Figure 5. The resulting occupancy grid has dimensions of occ$\in\{0,1\}^{1059\times 478\times 1059}$. To facilitate training, we divide the scene into smaller chunks of size $(50,h_{\text{vox}},50)$ along the $x,y,z$ axis, where $y$ represents the height axis. Chunks are sampled via occ[i-25:i+25,:,k-25:k+25], where $i$ and $k$ are sample coordinates along the $x$ and $z$ axis of the scene, respectively. The height of the sampled chunk $h_{\text{vox}}\leq 478$ is determined by the tallest occupied voxel within each chunk. As a result, chunk heights vary.
@@
Figure 6: We show generation results for the vector set and triplane baseline trained on a single scene. Orange boxes and zoomed-in renders highlight chunk outpainting coherence, while blue boxes illustrate scene geometry (zoom in for details). Zoomed-in renders may also have adjusted camera angles for clarity. The triplane model struggles with building details and introduces noisy artifacts, whereas the vector set model has better fidelity. While models can generate continuous chunks with smoothly connected roads and pipes (orange box, right), they sometimes fail to maintain coherence across chunks, leading to broken bridges (orange box, left).
@@ 3.2.1 Encoder
Our encoder $\mathcal{E}$ compresses sampled scene chunks $P$, which consists of varying heights into a uniform set of vectors. For each chunk $P_{i}$, we uniformly sample a fixed size point cloud $\bm{p}\in\mathbb{R}^{N_{p}\times 3}$ of $N_{p}$ points. Following 3DShape2Vecset [49], we use a cross-attention (CA) layer to aggregate the point clouds into a compact representation, leveraging a learnable set of fixed features with a reduced number of tokens. Next, a fully connected (FC) layer predicts the mean and variance, from which we apply the reparameterization trick [20] to generate embeddings for the chunks. This results in $\bm{z}^{\bm{p}}=\mathcal{E}(\bm{p})\in\mathbb{R}^{V\times c}$, where $V$ is the number of vectors and $c$ is the channel size.
However, we found that this is prone to posterior collapse [39], a common issue in VAEs. To address this, we introduce an alternative regularization strategy. Specifically, we sample another point cloud $\bm{q}\in\mathbb{R}^{N_{p}\times 3}$ from the same chunk $P_{i}$ and enforce consistency between their embeddings: $\mathcal{L}_{\text{emb}}=(\bm{z}^{\bm{p}}-\bm{z}^{\bm{q}})^{2}$. This prevents VAE collapse as it prevents the encoder from ignoring its input and enforces similar embeddings for point clouds sampled from the same chunk, while adding only a minimal overhead—just an additional pass through the CA and fully connected layers.
Furthermore, we predict the chunk’s height from its embedding to prune unnecessary occupancy predictions at inference time. Since the embedding may be generated by the diffusion model, we do not have access to the ground truth height of each chunk during inference. To account for this, we learn a height embedding $\bm{e}_{h}\in\mathbb{R}^{1\times c}$ and use it to query the latent representation, yielding the predicted height $\hat{h}=\text{FC}(\text{CA}(\bm{z^{p}},\bm{e}_{h}))$. The height loss is calculated as $\mathcal{L}_{\text{height}}=(\hat{h}-\tilde{h})^{2}$ where $\tilde{h}$ is the ground truth height.
@@ 3.2.2 Decoder
Next, we pass the embedding through $L$ self-attention (SA) layers in the decoder to obtain the output features $\bm{f}_{\text{out}}=\text{SA}^{(L)}\circ...\circ\text{SA}^{(1)}(\bm{z^{p}})$, which are then passed to either the vector set or triplane head to decode the occupancy.
Triplane. For the triplane head, $\bm{f}_{\text{out}}$ is reshaped into a triplane and upsampled using deconvolution layers, similar to LRM [15], yielding a feature grid $T\in\mathbb{R}^{3\times h_{\text{tri}}\times w_{\text{tri}}\times c_{\text{tri}}}$. Query points $\bm{r_{tri}}$ as outlined in Sec. 3.1 are normalized with $\bm{d}=(50,50*S,50)$, where the scale factor $S$ further compresses points along the y-axis. To ensure all coordinates remain within $[-1,1]$ for valid triplane sampling, we apply clamping. Note that we define $-1$ as the left or bottom bound of the triplane and $1$ as the right or top bound. Features are extracted via bilinear interpolation from the triplanes, concatenated, and processed by a fully connected layer to predict occupancies $\hat{o}_{r}$.
@@
Loss. We supervise the occupancy with binary cross entropy $\mathcal{L}_{\text{ce}}=\text{BCE}(\hat{o}_{r},\tilde{o}_{r})$ with the ground truth occupancy $\tilde{o}_{r}$. The embedding is supervised with a KL loss $\mathcal{L}_{\text{kl}}$. See Zhang et al. [49] for more details. The total loss is $\mathcal{L}=\lambda_{\text{kl}}\mathcal{L}_{\text{kl}}+\lambda_{\text{emb}}\mathcal{L}_{\text{emb}}+\lambda_{\text{ce}}\mathcal{L}_{\text{ce}}+\lambda_{\text{height}}\mathcal{L}_{\text{height}}$, and $\lambda$s are the loss weights. We note that, unlike the triplane method, the vector set allows query points to have ranges beyond $[-1,1]$. This is because cross-attention and positional encoding can directly handle arbitrary values without requiring them to be within a specific range for valid sampling, unlike spatial latent-based approaches like triplanes.
Inference Time Decoding. During inference, since we have no access to the encoding point cloud or ground truth height $\tilde{h}$, we use the predicted height $\hat{h}$ to prune unnecessary calculations. First, our model predicts $\hat{h}$. Then, we generate all query coordinates $\bm{r}_{\text{vox}}$ within a volume of shape $(50,50*(\hat{h}+1)/2,50)$, where each query point $\bm{r}$ is normalized via $12*(\bm{r}_{\text{vox}}/\bm{d})-$ 1. These normalized coordinates query the latent representation for occupancy predictions, and we apply marching cubes to the occupancy volume to reconstruct the chunk’s mesh.
@@ 3.3 Diffusion
Unlike existing models [26, 44, 32] that utilize RePaint [31] for outpainting, which requires extra diffusion steps, we propose a more efficient approach. Our outpainting diffusion model generates four chunks at once in a $2\times 2$ grid, conditioning on previously generated chunks using different patterns to handle all scenarios as shown in Figure 4 for continuous generation in a raster scan order.
We adopt DDPM [14] for diffusion probabilistic modeling. For diffusion training, we sample a quadrant of neighboring chunk latents $\{\bm{z}^{0},\bm{z}^{1},\bm{z}^{2},\bm{z}^{3}\}$ arranged in a $2\times 2$ grid sampled from the scene and encoded by the VAE. The order of chunks can be seen in Figure 4. At each training step we sample a Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ at a randomly sampled timestep $t\in[1,T]$ which is added to the quadrant of embeddings to obtain noisy latents $\bm{X}_{t}=\{\bm{x}^{0}_{t},\bm{x}^{1}_{t},\bm{x}^{2}_{t},\bm{x}^{3}_{t}\}\in\mathbb{R}^{4\times V\times c}$, where $\bm{x}^{i}_{t}\in\mathbb{R}^{V\times c}$ is the noised $i$-th latent at timestep $t$.
Our diffusion model takes noisy latents $\bm{X}_{t}$ as input, along with a binary mask indicating existing chunks and corresponding conditional embeddings providing context for outpainting. We consider four possible conditioning configurations, each with a corresponding mask $M$ and condition embeddings $Z_{\text{cond}}$, as illustrated in Figure 4. The mask is defined as $M=\{\bm{m}^{0},\bm{m}^{1},\bm{m}^{2},\bm{m}^{3}\}\in\mathbb{R}^{4\times V\times 1}$, where $M$ has four possible values: $\{\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0}\}$, $\{\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{0}\}$, $\{\mathbf{1},\mathbf{1},\mathbf{0},\mathbf{0}\}$ or $\{\mathbf{1},\mathbf{1},\mathbf{1},\mathbf{0}\}$, with $\mathbf{0},\mathbf{1}\in\mathbb{R}^{V\times 1}$. The condition embeddings are constructed based on the corresponding masks $\{\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0}\}$, $\{\bm{z}^{0},\mathbf{0},\bm{z}^{2},\mathbf{0}\}$, $\{\bm{z}^{0},\bm{z}^{1},\mathbf{0},\mathbf{0}\}$ and $\{\bm{z}^{0},\bm{z}^{1},\bm{z}^{2},\mathbf{0}\}$ with $\mathbf{0}\in\mathbb{R}^{V\times c}$ to form $Z_{\text{cond}}$. The mask, conditional embeddings and positional embeddings PE are concatenated to obtain the condition of the diffusion model:
@@
Note that during diffusion model training, one of the four configurations is randomly selected, along with the corresponding mask and conditional embeddings. Please see the supplemental (Algorithm 1) for more details on how the raster scan generation works during inference.

### v2

@@ Abstract
Figure 2: A scene generated by our model, trained on multiple scenes. We texture the scene using SceneTex [6]. As shown, our model combines elements from different scenes in the dataset such as castles and skyscrapers.
@@ 1 Introduction
Following the trend of diffusion models in image generation—such as Stable Diffusion [38]—generative models for creating standalone 3D objects [48, 24] are also rapidly maturing. The success of such models is partly driven by large-scale datasets like Objaverse [10]. However, large-scale scene generation remains limited, with many works focused on indoor environments [18, 47, 33, 4]. While some methods utilize semantic car datasets [26, 31, 27] or CAD models [30], these datasets lack high-quality geometry and have limited scene variation and height. Compared to previous methods, we tackle a difficult task with three key challenges: 1) enabling fast and efficient generation of outdoor scenes with varying heights, 2) generating landscapes with varying styles into a cohesive and continuous scene, 3) curating a dataset that facilitates exploration into such methods and supports joint training of heterogeneous scenes.
@@
Previous works on generating unbounded indoor scenes using 3D-Front [12] assume scenes can be divided into equal-sized cubes, and learn spatially structured latents like triplanes [47] or dense feature grids [33] for compression. This approach struggles with outdoor scenes featuring buildings like skyscrapers with significantly different heights. As shown in Figure˜3, naively scaling these representations can result in intensive memory usage. Rescaling the scene via normalization will also degrade details. To address this, we compress scene chunks into vector sets [52], achieving better compression than triplanes for more efficient training and higher quality generations. Additionally, we train our diffusion model for explicit outpainting, further improving inference speed over methods [47, 33] utilizing resampling-based inpainting [32]. Furthermore, prior methods focus on uniform datasets, such as indoor or urban driving scenes. In contrast, we aim to explore generative models that can bring together scenes with different context such as castles and cities as demonstrated by our model in Figure˜2.
Another key limitation is the lack of openly available high quality outdoor scenes for training. 3D-Front [12] is used by many prior works to train unbounded indoor scene generation. However, it lacks height diversity and non-furniture geometry. Semantic driving datasets [45, 2] and urban reconstructions [37, 14, 13] offer limited mesh quality. Objaverse [10] includes many scenes that could facilitate training for this task. However, several challenges remain. Notably, these meshes lack a unified global scale, and feature highly varied ground geometries—ranging from very thick to thin planes—making it difficult to train a generative model effectively as a homogeneous dataset. To address this, we curate a small set of scenes from Objaverse, clean ground geometries and establish a unified scale, enabling the development of methods that handle (1) diverse scene height distributions and (2) outdoor geometry from varied sources and styles. We show that our dataset pipeline effectively enables joint training of heterogeneous scenes in a unified model.
@@ 2 Related Work
Figure 4: Overview of our VAE and diffusion models. For the VAE we use vecset [52] for latent compression. For the diffusion model we train a conditional outpainting model with four different settings to allow for fast generation in a raster scan pattern during inference.
@@
3D Bounded Single Exemplar Scene Gen. Several works focus on generating 3D scenes from a single exemplar [44, 19, 46, 28], follow a coarse-to-fine progressive approach by capturing patch statistics across scales. Similar to SinGAN [39], SinGRAV [44], 3INGAN [19], and Wu and Zheng [46] adopt a pyramid of GANs, where discriminators operate on smaller patch sizes to enforce local consistency while allowing diverse global compositions. Li et al. [28] introduces EM-like optimization for patch matching and blending across different scales. Although these methods generate scenes with varying compositions and aspect ratios, they are limited to bounded generation, and producing an entire scene at once is memory intensive for larger scenes.
3D Unbounded Scene Gen. One common framework for 3D unbounded scene generation starts by dividing scenes into smaller chunks and using an autoencoder to compress them into latent representations. A diffusion model is then trained on the scene chunk latents. By overlapping the current chunk with previously generated ones, these methods can enable continuous outpainting by using resampling-based inpainting methods like RePaint [32].
@@
For unbounded indoor scene generation, methods typically train on 3D-Front [12] and assume cubic scene chunks. BlockFusion [47], like SemCity, compresses scene chunks into triplanes. It then uses another VAE to further compress the triplanes into a lower resolution, reducing redundancy in the representation to improve diffusion learning. LT3SD [33] encodes TUDF chunks into multi-level dense feature grid latent trees, organizing both its autoencoder and diffusion models in a hierarchical manner for coarse-to-fine chunk generation. While these representations perform well for indoor scenes with fixed ceiling heights, they struggle to scale to outdoor environments with significant height variations, as shown in Figure˜3. To tackle this, we use a vector set representation which compresses chunks into a uniform size, offering better compression and performance, and enable faster generation with a outpainting model.
Other methods take a different approach to unbounded outdoor generation. CityDreamer [49] and InfiniCity [30] leverage top-down views for unbounded generation but are limited to cityscapes with low-resolution geometry. SceneDreamer [7] and Persistent Nature [5] are promising directions to generate unbounded 3D worlds from only 2D images. However, without explicit 3D supervision, the resulting geometry lacks detail.
@@ 3 Method
We represent a large outdoor scene as a collection of local scene chunks, where a VAE learns to compress individual chunks into vector sets, and a diffusion model generates neighboring chunks in a $2\times 2$ grid using the VAE’s learned representation. By combining both models, we enable continuous and unbounded outdoor scene synthesis. Our training framework (see Figure˜4) follows the Latent Diffusion Model (LDM)[38].
We outline the dataset curation process and the sampling of training chunks in Sec.˜3.1. To explore efficient scene generation, we experiment with triplane and vector set latents. The VAE backbone and the process of obtaining occupancies from these representations are described in Sec.˜3.2. For fast inference, we train an explicit outpainting diffusion model, detailed in Sec.˜3.3. Optionally, textures can be generated using SceneTex [6] as shown in Sec.˜3.4.
@@ 3.1 Dataset
Filtering and Preprocessing. We begin by filtering Objaverse [10] using multi-view embeddings from DuoduoCLIP [25] to select 43 high-quality scenes. Next, we establish a consistent scale by labeling scenes with relative scales. We then convert each scene into an occupancy grid with a resolution multiplied by its labeled relative scale and generate meshes using the marching cubes algorithm for point cloud sampling. To ensure uniform ground geometry across scenes, we ensure a uniform thickness below the ground level for each scene. For more details, see the supplement.
@@
Chunk Sampling. A sample scene processed through our curation pipeline is shown in Figure˜5. The resulting occupancy grid occ has dimensions of $1059\times 478\times 1059$}. To facilitate training, we divide the scene into smaller chunks of size $(50,h_{\text{vox}},50)$ along the $x,y,z$ axis, where $y$ represents the height axis. Chunks are sampled via occ[i-25:i+25,:,k-25:k+25], where $i$ and $k$ are sample coordinates along the $x$ and $z$ axis of the scene, respectively. The height of the sampled chunk $h_{\text{vox}}\leq 478$ is determined by the tallest occupied voxel within each chunk.
@@
Figure 6: We show generation results for the vector set and triplane baseline trained on a single scene. Orange boxes and zoomed-in renders highlight chunk outpainting coherence, while purple boxes illustrate scene geometry (zoom in for details). Zoomed-in renders may also have adjusted camera angles for clarity. The triplane model struggles with building details and introduces noisy artifacts, whereas the vector set model has better fidelity. While models can generate continuous chunks with smoothly connected roads and pipes (orange box, right), they sometimes fail to maintain coherence across chunks, leading to broken bridges (orange box, left).
@@ 3.2.1 Encoder
Our encoder $\mathcal{E}$ compresses sampled scene chunks $P$ consisting of varying heights into a uniform set of vectors. For each chunk $P_{i}$, we uniformly sample a fixed size point cloud $\bm{p}\in\mathbb{R}^{N_{p}\times 3}$ of $N_{p}$ points. Following 3DShape2Vecset [52], we use a cross-attention (CA) layer to aggregate the point clouds into a compact representation, leveraging a learnable set of fixed features with a reduced number of tokens. Next, a fully connected (FC) layer predicts the mean and variance, from which we apply the reparameterization trick [21] to generate embeddings for the chunks. This results in $\bm{z}^{\bm{p}}=\mathcal{E}(\bm{p})\in\mathbb{R}^{V\times c}$, where $V$ is the number of vectors and $c$ is the channel size.
However, we found that this is prone to posterior collapse [42], a common issue in VAEs. To address this, we introduce an alternative regularization strategy. Specifically, we sample another point cloud $\bm{q}\in\mathbb{R}^{N_{p}\times 3}$ from the same chunk $P_{i}$ and enforce consistency between their embeddings: $\mathcal{L}_{\text{emb}}=(\bm{z}^{\bm{p}}-\bm{z}^{\bm{q}})^{2}$. This prevents VAE collapse as it prevents the encoder from ignoring its input and enforces similar embeddings for point clouds sampled from the same chunk, with minimal overhead—one extra CA and FC pass.
Furthermore, we predict the chunk’s height from its embedding to prune unnecessary occupancy predictions at inference time. At inference time, the embedding is generated by the diffusion model and thus we do not have access to the ground truth height of each chunk. To account for this, we learn a height embedding $\bm{e}_{h}\in\mathbb{R}^{1\times c}$ and use it to query the latent representation, yielding the predicted height $\hat{h}=\text{FC}(\text{CA}(\bm{z^{p}},\bm{e}_{h}))$. The height loss is calculated as $\mathcal{L}_{\text{height}}=(\hat{h}-\tilde{h})^{2}$ where $\tilde{h}$ is the ground truth height.
@@ 3.2.2 Decoder
Next, we pass the embedding through $L$ self-attention (SA) layers in the decoder to obtain the output features $\bm{f}_{\text{out}}=\text{SA}^{(L)}\circ\cdots\circ\text{SA}^{(1)}(\bm{z^{p}})$, which are then passed to either the vector set or triplane head to decode the occupancy.
Triplane. For the triplane head, $\bm{f}_{\text{out}}$ is reshaped into a triplane and upsampled via deconvolutions, following LRM [16], yielding a feature grid $T\in\mathbb{R}^{3\times h_{\text{tri}}\times w_{\text{tri}}\times c_{\text{tri}}}$. Query points $\bm{r_{tri}}$ as outlined in Sec.˜3.1 are normalized with $\bm{d}=(50,50*S,50)$, where the scale factor $S$ further compresses points along the y-axis. To ensure all coordinates remain within [1,1][-1,1] for valid triplane sampling, we apply clamping. Note that we define 1-1 as the left or bottom bound of the triplane and $1$ as the right or top bound. Features are extracted via bilinear interpolation from the triplanes, concatenated, and processed by a FC layer to predict occupancies $\hat{o}_{r}$.
@@
[…]
Inference Time Decoding. During inference, since we have no access to the ground truth height $\tilde{h}$, we use the predicted height $\hat{h}$ to prune unnecessary calculations. First, our model predicts $\hat{h}$. Then, we generate all query coordinates $\bm{r}_{\text{vox}}$ within a volume of shape $(50,50*(\hat{h}+1)/2,50)$, where each query point $\bm{r}$ is normalized via $2*(\bm{r}_{\text{vox}}/\bm{d})-1$. These normalized coordinates query the latent representation for occupancy predictions, and we apply marching cubes to the occupancy volume to reconstruct the chunk’s mesh.
@@ 3.3 Diffusion
Unlike existing models [27, 47, 33] that use RePaint [32] for outpainting, requiring extra diffusion steps, we propose a more efficient approach. Our outpainting diffusion model generates four chunks at once in a $2\times 2$ grid, conditioning on previously generated chunks using different patterns to handle all scenarios as shown in Figure˜4 for continuous generation in a raster scan order.
We adopt DDPM [15] for diffusion probabilistic modeling. During training, we sample a quadrant of neighboring chunk latents $\{\bm{z}^{0},\bm{z}^{1},\bm{z}^{2},\bm{z}^{3}\}$ arranged in a $2\times 2$ grid sampled from the scene and encoded by the VAE. The order of chunks is shown in Figure˜4. At each training step we sample a Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ at a randomly sampled timestep $t\in[1,T]$ which is added to the quadrant of embeddings to obtain noisy latents $\bm{X}_{t}=\{\bm{x}^{0}_{t},\bm{x}^{1}_{t},\bm{x}^{2}_{t},\bm{x}^{3}_{t}\}\in\mathbb{R}^{4\times V\times c}$, where $\bm{x}^{i}_{t}\in\mathbb{R}^{V\times c}$ is the noised $i$-th latent at timestep $t$.
Our diffusion model takes noisy latents $\bm{X}_{t}$ as input, along with a binary mask indicating existing chunks and corresponding conditional embeddings providing context for outpainting. We consider four possible conditioning configurations, each with a corresponding mask $M$ and condition embeddings $Z_{\text{cond}}$, as illustrated in Figure˜4. The mask is defined as $M=\{\bm{m}^{0},\bm{m}^{1},\bm{m}^{2},\bm{m}^{3}\}\in\mathbb{R}^{4\times V\times 1}$, where $M$ has four possible values: $\{\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0}\}$, $\{\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{0}\}$, $\{\mathbf{1},\mathbf{1},\mathbf{0},\mathbf{0}\}$ or $\{\mathbf{1},\mathbf{1},\mathbf{1},\mathbf{0}\}$, with $\mathbf{0},\mathbf{1}\in\mathbb{R}^{V\times 1}$. The condition embeddings are constructed based on the corresponding masks $\{\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0}\}$, $\{\bm{z}^{0},\mathbf{0},\bm{z}^{2},\mathbf{0}\}$, $\{\bm{z}^{0},\bm{z}^{1},\mathbf{0},\mathbf{0}\}$ and $\{\bm{z}^{0},\bm{z}^{1},\bm{z}^{2},\mathbf{0}\}$ with $\mathbf{0}\in\mathbb{R}^{V\times c}$ to form $Z_{\text{cond}}$. The mask, conditional embeddings and positional embeddings PE are concatenated to obtain the condition of the diffusion model:
@@
During diffusion model training, one of the four configurations is randomly selected, along with the corresponding mask and conditional embeddings. Please see the supplemental for more details on the raster scan generation.
3.4 Scene Texturing
We optionally texture generated scenes using SceneTex [6], modifying camera settings for better coverage of our scenes compared to the default Spherical Camera [^1] https://github.com/daveredrum/SceneTex?tab=readme-ov-file#cameras. Using Blender, we manually keyframe the scene with a fixed camera trajectory and speed, yielding 1.2k to 1.8k keyframes depending on scene shape and scale to use for SceneTex training. Scenes are normalized to $1,1]^{3}$, and we set the camera’s field of view to 40° with a near clipping distance of 0.01m. Please refer to the supplement for more details.